# Generated Contents Enrichment

### v2

@@ IV-D2 Generated Image Quality Metrics
TABLE II: Inception Scores (IS) [105] and Fréchet Inception Distances (FID) [106] are computed to compare real-world VG dataset and sets of images generated from GT scene descriptions (GT Synthesized), Simple descriptions, and Enriched versions utilizing Our model and ChatGPT with three different prompt styles.
Inception Scores (IS) [105] and Fréchet Inception Distances (FID) [106] are widely recognized metrics used to evaluate the quality of synthesized images. IS relies on a pre-trained image classifier to assess the presence of recognizable objects and their diversity within the generated image set. FID, on the other hand, employs a pre-trained classifier to extract deep features from both generated and real image sets. It then measures the statistical similarity between the distributions of these extracted features for the generated and real image sets.
Table II exhibits FID and IS metrics for evaluating the image quality. Stable Diffusion is employed as the image synthesizer to generate images for different scenarios compared to the real-world VG images:
  • First, simple scene descriptions from the test set are utilized for synthesizing a total of 36.5k simple images denoted by Simple.
  • Second, the same simple descriptions are fed to our proposed GCE method or ChatGPT with the three introduced prompting styles. This is to produce the enriched descriptions used for synthesizing 36.5k enriched images denoted by Enriched for each of the mentioned four cases: Enriched ChatGPT Direct/Scene/Object and Enriched Ours.
  • Third, GT training scene descriptions are directly used to generate 35k images denoted by GT Synthesized.
  • The fourth set of images is 87k real-world VG images denoted by VG.
@@
Objs Acc 19.04 18.33 18.52 18.66 17.88 19.48
Avail Preds Acc 41.95 37.43 38.87 37.47 36.08 42.20
Not Avail Preds Acc - - - - - -
Avail Edges Acc 75.00 74.89 74.20 75.33 74.62 75.18
Not Avail Edges Acc 99.58 99.61 99.61 99.60 99.66 99.57
Scene Class Acc 20.33 19.99 21.14 21.66 - 22.66
TABLE III: Ablation study results. Modules are altered with respect to SG Enrich as the baseline.
IV-E Ablation Study
In order to demonstrate the necessity of each component in our model and their individual contribution to the overall performance, we conducted an ablation study. The results of this study are presented in Table III, where the ablated models are compared employing our defined metrics for scene graph enrichment. Since this task lacks a unique solution, raw results with the specified metrics may be misleading for comparison and often can contradict human judgments.
@@ IV-E Ablation Study
IV-F User Studies
@@
Stable Diffusion Image Real-World Image $\pm$ Standard Deviation
@@
Enriched Image Simple Image $\pm$ Standard Deviation
@@
Enriched Image Simple Image $\pm$ Standard Deviation
@@
TABLE IV: User studies are conducted in four sections to evaluate the effectiveness of the enrichment process.
Our efforts to demonstrate the performance of our model and its different variations involved employing our defined metrics and borrowed measures from the assessment of generated image quality. Nevertheless, it is crucial to acknowledge that evaluating the model’s performance solely against GT provides only a rough measure due to the inherently ill-posed nature of the problem. Ultimately, human judgment is essential to truly claim the quality of the enrichment process. In this regard, our user studies, detailed in Table IV, were conducted with the participation of 110 individuals, spanning four sections.
S1: Real vs. Stable Diffusion Images Realisticness. To showcase the capability of our image generator to produce realistic images, we presented users with ten pairs of images. Each pair consists of an image selected from either the VG or MS COCO [107] and an image generated by Stable Diffusion, both depicting similar scenes. Participants were asked: Which one is more realistic? Surprisingly, only in 20.09% of cases did participants select the synthesized images as more realistic. In most instances, users tend to successfully identify the generated images, probably attributed to the vivid and visually appealing nature of Stable Diffusion images, imitating the characteristics of their training data.
@@ V Conclusion
A Preds Loss Weight 100 50 100 50
NA Preds Loss Weight 0.1 0.01 5 0.01
Edges Loss Weight 1 0.01 100 10
@@ D Update Every 200 1 - 1
TABLE VI: Quantitative Results, including our defined measures for the scene graph enrichment task, are exhibited for selected hyperparameter settings.
@@
To show the iterative power of the enrichment process, the model is forced to enrich an object not already present in the input SG at each step during inference. Adam optimizer [108] is used with $10^{-4}$ as the learning rate. Also, We employed early termination for training our network based on the accuracies of the validation split. Using the proposed model on the VG dataset, continuing the training leads to improvement in training metrics and worsening validation metrics, indicating overfitting.
@@
Table V describes the various settings used for hyperparameter tuning. Before training the model end-to-end, a hyperparameter optimization is performed, involving only our SG generator and discriminators. In other words, no image synthesizer is used at that phase. That is because training and inference time are increased while disconnecting the two mentioned components. Accordingly, the underlined values in Table V are indeed eliminated from the final hyperparameter tuning for the end-to-end model, as they performed poorly in the first stage. As already mentioned, loss weights have a crucial impact on the performance of our model. For instance, assigning a higher weight to Not Avail Preds could result in a model that outputs an enriching object without any enriching edges.
@@
Hyperparameter tuning is performed as a sweep of different values described in Table V. Best models are picked based on a better summation of validation accuracies in the model. Among them, the ones that performed qualitatively better on a limited selection of SGs are chosen. As so many possible combinations of hyperparameters result in a huge space to discover, we select a limited combination of the parameters based on Bayes hyperparameter optimization as we train around 650 different instances.
@@
The batch size varies from 32 to 128 scene graphs. A too-low value for batch size leads to the GPU memory and processing power not being fully utilized. A too-high value for it results in longer training time as the GT values for each SG direct the weights updates in so many different directions. This may result in a poor update at each mini-batch iteration. Dropout layers are not used in the GCN version of [3], as the authors reported that using them reduces performance. Our experiments also show the same result for higher probability values of dropout. However, we found some lower values to be beneficial for overfitting prevention. Normalization layers are not used in [3], but they potentially could aid the model to converge faster. As a result, we tried out Batch Normalization [109], Layer Normalization [110], or no normalization layer at all for our graph convolutions.
@@
Too small values for the object loss weight or available predicates loss weight are eliminated based on our preliminary experiments, as these are two of the fundamentals of our overall objective. The option to set a weight of zero to Not Avail Preds loss is available to give the network the opportunity to update its weights based on only available predicates. In other words, when there is no edge between two objects in a scene graph, that does not mean the only possible reality is that these two do not have a relationship in a scene. Suppose the Enriching Edge Detector selects an edge where the predicate is unavailable in the GT. In that case, we may force the Enriching Predicate GCN to classify that edge as $none\_pred$, indicating no relationship based on the GT or not penalize the output predicate as the GT is indeed unavailable. This is controlled using different values for predicate loss weights.
@@
The embedded dimension $D_{emb}$ for the generator is indeed the first GConv’s $D_{in}$ selected from 16 to 256. $x_{1}\>x_{2}\>...\>x_{n}$ architecture for GCN means there are $n-1$ GConv components stacked together where the $i^{th}$ GConv’s input, output, and hidden dimensions are $x_{i}D_{emb}$, $x_{i+1}D_{emb}$, and $2(x_{i}+x_{i+1})D_{emb}$, respectively.

### v3

@@ 4.4.2 Generated Image Quality Metrics
Inception Scores (IS) (Salimans et al., 2016) and Fréchet Inception Distances (FID) (Heusel et al., 2017) are widely recognized metrics used to evaluate the quality of synthesized images. IS relies on a pre-trained image classifier to assess the presence of recognizable objects and their diversity within the generated image set. FID, on the other hand, employs a pre-trained classifier to extract deep features from both generated and real image sets. It then measures the statistical similarity between the distributions of these extracted features for the generated and real image sets.
Table 2 exhibits FID and IS metrics for evaluating the image quality. Stable Diffusion is employed as the image synthesizer to generate images for different scenarios compared to the real-world VG images: (i) First, simple scene descriptions from the test set are utilized for synthesizing a total of 36.5k simple images denoted by Simple. These represent synthesized images produced by a SOTA image generator (Rombach et al., 2022) without incorporating any explicit content enrichment model. (ii) Second, the same simple descriptions are fed to our proposed GCE method or ChatGPT with the three introduced prompting styles. This is to produce the enriched descriptions used for synthesizing 36.5k enriched images denoted by Enriched for each of the mentioned four cases: Enriched ChatGPT Direct/Scene/Object and Enriched Ours. (iii) Third, GT training scene descriptions are directly used to generate 35k images denoted by GT Synthesized. (iv) The fourth set of images is 87k real-world VG images denoted by VG.
@@
Table 3: Ablation study results. Modules are altered with respect to SG Enrich as the baseline.
@@
Obj. Acc. 19.04 18.33 18.52 18.66 17.88 19.48
Avail. Pred. Acc. 41.95 37.43 38.87 37.47 36.08 42.20
Not Avail. Pred. Acc. - - - - - -
Avail. Edge Acc. 75.00 74.89 74.20 75.33 74.62 75.18
Not Avail. Edge Acc. 99.58 99.61 99.61 99.60 99.66 99.57
Scene Class. Acc. 20.33 19.99 21.14 21.66 - 22.66
4.5 Ablation Study
In order to demonstrate the necessity of each component in our model and their individual contribution to the overall performance, we conducted an ablation study. The results of this study are presented in Table 3, where the ablated models are compared employing our defined metrics for scene graph enrichment. Since this task lacks a unique solution, raw results with the specified metrics may be misleading for comparison and often can contradict human judgments.
@@ 4.5 Ablation Study
4.6 User Studies
Table 4: User studies with 110 participants are conducted to assess the performance and realism of generated images. S1 shows that users can generally distinguish AI-generated images, probably due to their vivid appearance. S2 highlights that our enriched images are perceived as realistic and structurally plausible. S3 reveals our model generates enriched content while reflecting the original description. S4 confirms user preference for our enriched images due to their coherence and meaningful content.
@@ 4.6 User Studies
(A) Stable Diffusion Image (B) Real-World Image $\pm$ Standard Deviation
@@
(A) Enriched Image (B) Simple Image $\pm$ Standard Deviation
@@
(A) Enriched Image (B) Simple Image $\pm$ Standard Deviation
@@
Our efforts to demonstrate the performance of our model and its different variations involved employing our defined metrics and borrowed measures from the assessment of generated image quality. Nevertheless, it is crucial to acknowledge that evaluating the model’s performance solely against GT provides only a rough measure due to the inherently ill-posed nature of the problem. Ultimately, human judgment is essential to truly claim the quality of the enrichment process. In this regard, our user studies, detailed in Table 4, were conducted with the participation of 110 individuals, spanning four sections.
S1: Real vs. Stable Diffusion Images Realisticness. To showcase the capability of the image generator to produce realistic images, we presented users with ten pairs of images. Each pair consists of an image selected from either the VG or MS COCO (Lin et al., 2014) and an image generated by Stable Diffusion, both depicting similar scenes. Participants were asked: Which one is more realistic? Surprisingly, only in 20.09% of cases did participants select the synthesized images as more realistic. In most instances, users tend to successfully identify the AI-generated images, probably attributed to the vivid and visually appealing nature of Stable Diffusion images, which imitate the characteristics of their training data.
@@ 5 Conclusion
Appendix A Method
A.1 GConv Operation
Our GCN and its building block GConv are depicted in Fig. 4. Elements of an edge are $(o_{i},r,o_{j})$ with corresponding vector values $(v_{i},v_{r},v_{j})$ and dimensions $(D_{in},D_{in},D_{in})$. They are concatenated and then fed into a graph convolution layer. At the start of the GConv operation, there are two fully connected (FC) layers with output dimensions $H=2(D_{in}+D_{out})$ and $3H$. The output of these FC layers is then divided into three parts, which represent candidate vectors for the subject node, the edge, and the object node. They are represented as $(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})$ with dimensions $(H,H,H)$. $\bar{v}_{r}$ is then fed into additional layers for the skip connection. If this operation is performed for every edge, in the end, for each node, we may have some candidate object vectors and some other candidate subject vectors. By computing the mean of these candidate vectors for each node and then feeding into additional FC layers with output dimensions $H$ and $D_{out}$, a new vector is calculated for each node after the skip connection. A new graph could be formed with the new vectors for the edges and nodes, which is then fed into the following graph convolution layers. There is also an additional skip connection link from the input graph to the final layers of the GCN. This is because we may train deeper networks using skip connections without gradient vanishing problems.
A.2 GCN Implementation
For processing a mini-batch in parallel, all the edges in the mini-batch graphs are fed into a graph convolution layer simultaneously. Similar to Johnson et al. (2018), in order to make each SG separate from others, it is made connected using a dummy node with $image$ as its category. Each node of the SG is connected to this dummy node with directed edges and $in\_image$ as their categories. In addition, two other vectors corresponding to objects and edges showing the corresponding SG placement in the mini-batch are fed to the networks.
Figure 4: Graph Convolutional Network (GCN) and its building block GConv are employed in Stage 1: Generative Adversarial SG Enrichment.
Figure 5: Enriching Edge Detector receives the nodes of a graph but not their edges. It also accepts hidden object vectors from the Enriching Object GCN as input. After concatenation, each node is fed to two neural networks with the same architecture but separate sets of weights to form hidden vectors for subjects and objects of a relationship. This part aims to transfer the nodes to another space where cosine similarity represents the potential of existing edges between two nodes. Therefore, the resulting subject and object vectors form a score matrix implemented as matrix multiplication.
A.3 Generator’s GCNs Architecture
The architecture of the generator’s two GCNs is the same but with different weights. First, the dimension of the data is increased to form a subspace where we can do the processing, then reduced back to a low dimension to encourage the network to extract an encoded version of input information. Finally, it returns to its original dimension. The number of neural network layers in the GCN classifiers and Enriching Edge Detector model are selected via GCN Classifier Layers and Enriching Edge Layers in Table 5. In addition, at the first layers of our GCNs, two embedding linear layers without bias work as object and predicate encoders to transform the one-hot-encoded categories of objects and relationships into another continuous space.
A.4 CLIP Text Encoder
As the CLIP text encoder cannot handle the scene graphs directly, each edge is converted to a sentence, including the subject, the predicate, and the object of the relationship. These sentences are fed to the CLIP text encoder in addition to the objects that do not participate in any relationship as a text prompt.
Appendix B Experiments
B.1 Image Generator
In some cases, scenes and objects synthesized with SG2IM (Johnson et al., 2018) are unrecognizable even for humans, which impacts the understanding of the scene by the Stage3: Visual Evaluation & Alignment. We attempted to utilize other SOTA image generators instead of SG2IM during training, such as Stable Diffusion (Rombach et al., 2022), which produces $512\times 512$ and $768\times 768$ images in the most recent version. However, because of the iterative nature of their inference, generating every single image of the mini-batch without considering the backpropagation is computationally expensive. Therefore, it is impossible to get any reasonable results while doing the hyperparameter tuning in a timely manner employing more sophisticated diffusion-based high-resolution image generators.
B.2 Pair of Discriminators
On the VG dataset with the selected architectures, distinguishing the enriched SGs from the original data is usually far more manageable than the enrichment task. That is why our discriminators often perform far better than the generator. We tried some instances to limit the discriminator compared to the generator counterpart by having an embedded dimension selected from 16 to 256, smaller than the generator’s embedded dimension. Updating the discriminators after each generator update makes them far more powerful than their counterpart generator. That is why different settings for updating the discriminators only one time for several updates of the generator are utilized based on Update Every in Table 5.
B.3 Training Details & Hyperparameter Tuning
Table 5: Hyperparameter Selections for training the network. As several loss terms are involved in the training process and considering that different values for their weights significantly impact the overall performance, we did a sweep over these options. Also, for training speed and generalizability, we tried out different options for dropout, normalization, and architecture settings. Besides, GConv HP Category is applied in both the generator and the discriminators. Also, the embedded dimension determines the data dimension after feeding the vectors to the embedding layers. In addition, the underlined values are eliminated using preliminary training only on the Generative Adversarial Scene Graph Enrichment module.
HP Category HP Values
Batch Size 32, 64, 128
GConv (D, G) Dropout 0, 0.05, 0.1, 0.15, 0.25, 0.5
Normalization none, batch, layer
Activation ReLU, LeakyReLU, PReLU
Loss Obj. Weight 0.001, 0.01, 0.1, 0.5, 1, 5, 10, 50, 100, 200, 500, 1000
Avail. Pred Weight 0.001, 0.01, 0.1, 0.5, 1, 5, 10, 50, 100
Not Avail. Pred Weight 0, 0.001, 0.01, 0.1, 0.5, 1, 5, 10, 50, 100
Edge Weight 0.001, 0.01, 0.1, 0.5, 1, 5, 10, 50, 100
GAN Weight 0, 0.001, 0.01, 0.1, 0.5, 1, 5, 10
Scene Features Weight 0.001, 0.01, 0.1, 0.5, 1, 5, 10, 50, 100, 200, 500, 1000
[…]
Image-SG Alignment Weight 1, 10, 100, 1000, 2000, 5000, 10000
Generator Embedded Dim. 16, 32, 64, 128, 256
GCN Architecture (a) 1 1 1 1 $\>1\>1\>1\>$ 1 1 1
(b) 1 1 1 1 1 1 $\>1\>1\>1\>1\>$ 1 1 1 1
(c) 1 4 2 2 4 11422411\>4\>2\>2\>4\>11 4 2 2 4 1
(d) 1 4 2 1 1 2 4 1142112411\>4\>2\>1\>1\>2\>4\>11 4 2 1 1 2 4 1
(e) 1 4 2 11212 1 2 4 $11\>4\>2\>1\>\frac{1}{2}\>\frac{1}{2}\>1\>2\>4\>$ 1
(f) 1 4 2 112141412 1 2 4 $11\>4\>2\>1\>\frac{1}{2}\>\frac{1}{4}\>\frac{1}{4}\>\frac{1}{2}\>1\>2\>4\>$ 1
FC Dropout 0.1, 0.25, 0.5
FC Norm none, batch, layer
FC Activation ReLU, LeakyReLU, PReLU
GCN Classifier Layers 1, 2, 4
Enriching Edge Layers 2, 4
Discriminator Embedded Dim. 16, 32, 64, 128, 256
GCN Architecture (a) $1\>\frac{1}{2}\>\frac{1}{4}\>\frac{1}{8}\>\frac{1}{8}\>\frac{1}{8}$
(b) 1 $1\>1\>\frac{1}{2}\>\frac{1}{2}\>\frac{1}{4}\>\frac{1}{4}\>\frac{1}{8}\>\frac{1}{8}\>\frac{1}{8}$
(c) 1 2 2 1 $1\>2\>2\>1\>1\>\frac{1}{2}\>\frac{1}{2}\>\frac{1}{4}\>\frac{1}{4}\>\frac{1}{8}\>\frac{1}{8}\>\frac{1}{8}$
Update Every 1, 5, 10, 50, 100, 200, 500, 1000, 2000
Table 6: Quantitative Results, including our defined measures for the scene graph enrichment task, are exhibited for selected hyperparameter settings.
Measure/HP SG Enrich Deep G Best NoD 128 BS
Obj. Acc. 19.04 17.31 15.09 16.85
Avail. Pred. Acc. 41.95 39.90 40.25 31.38
Not Avail. Pred. Acc. - - 3.02 -
Avail. Edge Acc. 75.00 75.64 76.08 66.89
Not Avail. Edge Acc. 99.58 99.70 99.54 99.54
Scene Class. Acc. 20.33 20.84 19.18 25.33
Mini-Batch Size 32 32 32 128
GConv Dropout 0.1 0.05 0.05 0
GConv Normalization none none BN none
GConv Activation LeReLU LeReLU PReLU ReLU
Obj. Loss Weight 1000 500 1000 200
A. Pred Loss Weight 100 50 100 50
N.A. Pred Loss Weight 0.1 0.01 5 0.01
Edge Loss Weight 1 0.01 100 10
@@ D Update Every 200 1 - 1
To show the iterative power of the enrichment process, the model is forced to enrich an object not already present in the input SG at each step during inference. Adam optimizer (Kingma & Ba, 2014) is used with $10^{-4}$ as the learning rate. Also, We employed early termination for training our network based on the accuracies of the validation split. Using the proposed model on the VG dataset, continuing the training leads to improvement in training metrics and worsening validation metrics, indicating overfitting.
@@
Table 5 describes the various settings used for hyperparameter tuning. Before training the model end-to-end, a hyperparameter optimization is performed, involving only our SG generator and discriminators. In other words, no image synthesizer is used at that phase. That is because training and inference time are increased while disconnecting the two mentioned components. Accordingly, the underlined values in Table 5 are indeed eliminated from the final hyperparameter tuning for the end-to-end model, as they performed poorly in the first stage. As already mentioned, loss weights have a crucial impact on the performance of our model. For instance, assigning a higher weight to Not Avail. Pred. could result in a model that outputs an enriching object without any enriching edges.
@@
Hyperparameter tuning is performed as a sweep of different values described in Table 5. Best models are picked based on a better summation of validation accuracies in the model. Among them, the ones that performed qualitatively better on a limited selection of SGs are chosen. As so many possible combinations of hyperparameters result in a huge space to discover, we select a limited combination of the parameters based on Bayes hyperparameter optimization as we train around 650 different instances.
@@
The mini-batch size varies from 32 to 128 scene graphs. A too-low value for mini-batch size leads to the GPU memory and processing power not being fully utilized. A too-high value for it results in longer training time as the GT values for each SG direct the weights updates in so many different directions. This may result in a poor update at each mini-batch iteration. Dropout layers are not used in the GCN version of Johnson et al. (2018), as the authors reported that using them reduces performance. Our experiments also show the same result for higher probability values of dropout. However, we found some lower values to be beneficial for overfitting prevention. Normalization layers are not used in (Johnson et al., 2018), but they potentially could aid the model to converge faster. As a result, we tried out Batch Normalization (Ioffe & Szegedy, 2015), Layer Normalization (Ba et al., 2016), or no normalization layer at all for our graph convolutions.
@@
Too small values for the object loss weight or available predicates loss weight are eliminated based on our preliminary experiments, as these are two of the fundamentals of our overall objective. The option to set a weight of zero to Not Avail. Pred. loss is available to give the network the opportunity to update its weights based on only available predicates. In other words, when there is no edge between two objects in a scene graph, that does not mean the only possible reality is that these two do not have a relationship in a scene. Suppose the Enriching Edge Detector selects an edge where the predicate is unavailable in the GT. In that case, we may force the Enriching Predicate GCN to classify that edge as $none\_pred$, indicating no relationship based on the GT or not penalize the output predicate as the GT is indeed unavailable. This is controlled using different values for predicate loss weights.
@@
The embedded dimmension $D_{emb}$ for the generator is indeed the first GConv’s $D_{in}$ selected from 16 to 256. $x_{1}\>x_{2}\>...\>x_{n}$ architecture for GCN means there are $n-1$ GConv components stacked together where the $i^{th}$ GConv’s input, output, and hidden dimensions are $x_{i}D_{emb}$, $x_{i+1}D_{emb}$, and $2(x_{i}+x_{i+1})D_{emb}$, respectively.